OpenStudy (anonymous):

Can someone please help me with this... A stone is dropped from the edge of a roof, and hits the ground with a velocity of −185 feet per second. Assume the acceleration due to gravity is -32 feet per second squared. How high (in feet) is the roof?

OpenStudy (amistre64):

since the stone is dropped it has no initial velocity

OpenStudy (amistre64):

the acc dtg, is integrated upwards to velocity: v(t) = -32t + C; where of course C = 0 since there is no initial velocity v(t) = -32t; when v(t) = -185, t=?

OpenStudy (amistre64):

-185 = -32t -185/-23 = t 5.78125 = t, this gives us how long it took for the stone to hit the ground

OpenStudy (amistre64):

the integration of velocity gives us position: h(t) = -16t^2 + Vi t + Hi Vi = initial velocity, =0 and we want to solve for Hi, knowing that at t=5.78125, h(5.78125) = 0 h(t) = -16t^2 + Hi h(5.78125) = -16(5.78125)^2 + Hi 0 = -16(33.4228515625) + Hi 16(33.4228515625) = Hi 534.765625 = Hi, or initial Height in feet

OpenStudy (anonymous):

Thanks!

OpenStudy (amistre64):

yep, hope it made sense :)